### 16 Mar 2026 v1

# Aperiodic Structures Never Collapse: Fibonacci Hierarchies for Lossless Compression

Abstract
Hierarchical dictionary methods lose effectiveness at deep scales when their underlying structure collapses after finitely many levels. We prove an Aperiodic Hierarchy Advantage for Fibonacci quasicrystal tilings: unlike periodic tilings, the Fibonacci hierarchy never collapses, preserving non-zero $n$-gram lookup positions at every depth and maintaining scale-invariant reuse capacity. From this structure we derive constant potential word coverage $W\varphi/\sqrt{5}$, maximal Sturmian codebook coverage efficiency $C_{m}/(F_{m}{+}1)$, bounded parsing overhead $1/\varphi$ bits per word, and lower coding entropy than comparable periodic hierarchies on long-range-dependent sources.
Our analysis is supported by quantitative results on golden compensation, Sturmian minimality, entropy, and redundancy. The Golden Compensation theorem shows that exponential decay in position count is exactly balanced by exponential growth in phrase length, making reuse capacity invariant across scales. The Sturmian Codebook Efficiency theorem links the minimal complexity law p(n)=n+1p(n)=n{+}1 to maximal coverage efficiency, while the redundancy analysis shows super-exponential decay $O(e^{-\varphi^{m}/\lambda})$ for the Fibonacci hierarchy, in contrast to the finite-depth lock-in of periodic tilings.
We validate these results with Quasicryth, [^1] Quasicryth: portmanteau of quasicrystal and compression; the name also reflects the internal structure — quasicrystalline tiling hierarchy. a lossless text compressor built on a ten-level Fibonacci hierarchy with phrase lengths $\{2,3,5,8,13,21,34,55,89,144\}$ words. In controlled A/B experiments with identical codebooks, the aperiodic advantage over Period-5 grows from 1,372 B at 3 MB to 1,349,371 B at 1 GB, explained by the activation of deeper hierarchy levels. On enwik9 (1 GB), Quasicryth achieves 359,883,431 B (35.99%), with 45,608,715 B attributable to the quasicrystal tiling alone.
Keywords: lossless compression, quasicrystal tiling, Fibonacci substitution, aperiodic hierarchy, Sturmian sequences, arithmetic coding, n-gram codebook, Pisot-Vijayaraghavan numbers
1. Introduction
Shannon’s equivalence of compression and prediction [1] implies that superior structural models enable superior codes. Classical lossless compressors exploit structure at the byte level [20, 18, 19] or through Burrows–Wheeler block transformations [21]. Word-level compressors [27] apply frequency-ranked codebooks, trading byte granularity for semantic structure. In all cases, the parsing strategy — which sequence lengths to attempt at each position — is fixed or locally adaptive.
Quasicryth introduces a fundamentally different paradigm: the parsing strategy itself is determined by a one-dimensional quasicrystal. The Fibonacci tiling $\sigma:L\to LS,\;S\to L$ generates an aperiodic binary sequence in which the position and scale of every super-tile are geometrically determined, providing $n$-gram lookup positions at phrase lengths $\{2,3,5,8,13,21,34,55,89,144\}$ words simultaneously. The decisive property is that this hierarchy never collapses: unlike any periodic tiling, the Fibonacci quasicrystal sustains both tile types at every scale, because the golden ratio $\varphi$ is a Pisot-Vijayaraghavan number and the tiling is Sturmian.
This paper makes the following contributions:
  1. 1.
    Provable hierarchy non-collapse. A formal proof (Section 4) that the Fibonacci substitution hierarchy extends indefinitely, while all periodic tilings with period $p$ collapse within $\lceil\log_{\varphi}(p)\rceil$ levels. The proof uses the Perron-Frobenius eigenvector analysis of the substitution matrix, the PV property of $\varphi$, and Weyl’s equidistribution theorem [2].
  2. 2.
    Sturmian structure as compression engine. The Fibonacci word is the canonical Sturmian sequence [3]: aperiodic, 1-balanced, and of minimal factor complexity n+1n{+}1. We show how these three properties translate directly into compression benefits: balance ensures uniform codebook coverage, aperiodicity prevents hierarchy collapse, and minimal complexity maximises codebook entry reuse — formalised as the Sturmian Codebook Efficiency theorem (Theorem 7) and the Goldilocks corollary (Corollary 8).
  3. 3.
    Deep hierarchy with ten levels. Quasicryth v5.5 implements the full Fibonacci substitution hierarchy to depth $k=9$, providing lookup positions for phrases of up to 144 words. At enwik9 scale (298 M words), the 89-gram and 144-gram levels activate for the first time, contributing 2,396 and 945 deep hits respectively — positions entirely unavailable to any periodic tiling.
  4. 4.
    Aperiodic Hierarchy Advantage Theorem (Theorem 1). Our central result (§4.1) characterises the Fibonacci quasicrystal as the only binary tiling, within the class of infinite aperiodic tilings, that preserves dictionary reuse at all hierarchy depths simultaneously. Non-collapse, scale-invariant coverage, maximum codebook efficiency, bounded overhead, and a strict coding entropy advantage over any periodic alternative are all provable from first principles. Eight supporting theorems (§4.24.17) provide the quantitative details:
    • Sturmian Codebook Efficiency (Theorem 7): the Fibonacci tiling has exactly $F_{m}{+}1$ distinct tile-type patterns at level $m$, the minimum for any aperiodic sequence. A codebook of $C_{m}$ entries achieves coverage efficiency $C_{m}/(F_{m}{+}1)$, provably maximum among all aperiodic binary tilings of the same depth. The Fibonacci tiling is characterised as the only aperiodic sequence satisfying both non-collapse and maximum efficiency at every level (Corollary 8).
    • Golden Compensation (Theorem 9): the potential word coverage $C(m)=P(m)\cdot F_{m}\to W\varphi/\sqrt{5}$ is the same at every hierarchy level. All structural variation is carried by the codebook hit rate $r(m)$ alone; the positional and phrase-length factors cancel exactly via $\varphi^{2}=\varphi+1$ and Binet’s formula.
    • Activation Threshold (Theorem 12): hierarchy level $m$ first contributes when $W\geq W^{*}_{m}=T_{m}\varphi^{m-1}/r_{m}$, giving a closed-form prediction for the corpus scale at which each deep level activates.
[…]
    • Strict Coding Entropy (Theorem 16): for any source with long-range phrase dependencies ($m$-LRPD), the per-word coding entropy of the Fibonacci hierarchy is strictly lower than that of any periodic tiling collapsed before depth $m$ — a fully information-theoretic inequality, not just a combinatorial one. Natural language corpora are $m$-LRPD for $m$ corresponding to phrase lengths up to at least 144 words (Corollary 17).
    • Fibonacci Redundancy Bound (Theorem 18): for exponentially mixing sources, the coding redundancy at Fibonacci level $m$ decays super-exponentially as $O(e^{-\varphi^{m}/(\lambda\sqrt{5})})$, while any periodic tiling is locked at fixed redundancy $\Omega(e^{-F_{m^{*}}/\lambda})$. The redundancy ratio $\to 0$ as $m\to\infty$.
    • Exponential Dictionary Efficiency (Theorem 19): the per-entry compression gain $E_{m}=F_{m}\bar{h}-\log_{2}C_{m}$ grows as $\Omega(\varphi^{m})$; total dictionary value for Fibonacci grows exponentially in $k_{\max}$, versus a fixed constant for any periodic tiling.
    • Convergent Flag Overhead (Theorem 14): the per-word entropy of the hit/miss flag stream for the Fibonacci tiling is bounded above by $1/\varphi\approx 0.618$ bits/word, regardless of hierarchy depth. Combined with the unbounded growth of $h_{\text{saved}}$, this implies that the net per-word compression efficiency of the Fibonacci tiling exceeds that of any periodic alternative for all sufficiently large $W$ (Corollary 15).
  5. 5.
    Quantified aperiodic advantage at scale. Controlled A/B experiments: at 100 MB the Fibonacci tiling saves 40,385 B over Period-5; at 1 GB this grows to 1,349,371 B. The 89-gram and 144-gram levels activate at 1 GB (enwik9), contributing 2,396 and 945 deep hits — positions structurally unavailable to any periodic tiling.
  6. 6.
    Single-parameter tiling. The entire parsing structure — all tile boundaries, hierarchy levels, deep n-gram positions, and adaptive model assignments — is determined by a single 2-byte phase parameter. The decoder regenerates the complete tiling deterministically; no structural information appears in the compressed bitstream.
  7. 7.
    Separate bz2 escape stream. Out-of-vocabulary words are segregated into a parallel stream compressed with bzip2 [22], achieving $\sim$2.6 bpb versus $\sim$3.5 bpb for inline arithmetic coding. The quasicrystalline tiling governs which words escape, maintaining the structural integrity of both streams.
The remainder of the paper is organised as follows. Section 2 surveys related work. Section 3 describes the algorithm pipeline. Section 4 develops the formal theoretical analysis. Sections 56 present the experimental setup and results. Section 7 reports the A/B ablation study. Section 8 concludes with future directions.
2. Related Work
2.1 Classical Lossless Compression
Huffman coding [18] assigns variable-length binary codes by symbol frequency; arithmetic coding [19] removes the integer-bit constraint, approaching the entropy bound to within a fraction of a bit. LZ77 [20] and its derivatives (gzip/DEFLATE, LZMA/xz, Zstandard) find repeated substrings within a sliding window. Bzip2 [22] applies the Burrows-Wheeler transform [21] followed by move-to-front and Huffman coding, achieving strong compression on structured text. Prediction by Partial Matching (PPM) [23] builds high-order adaptive context models with arithmetic coding. These approaches operate at the byte level and exploit local repetitions. Quasicryth is a word-level compressor and exploits hierarchical phrase structure rather than byte-level redundancy.
2.2 Word-Level and Phrase-Level Compression
Word-based models [27, 28] replace byte alphabets with word vocabularies, achieving better compression on natural language by exploiting morphological and lexical structure. LZW-style online dictionaries build phrase codebooks adaptively [29]. Static frequency-ranked codebooks are used in specialized formats for natural language text. Quasicryth uses static multi-level codebooks built from the full input, but the parsing of the input into phrases is determined by the quasicrystalline tiling rather than by greedy matching or fixed-length frames.
2.3 Quasicrystals and Aperiodic Tilings
Quasicrystals were discovered experimentally in 1984 by Shechtman et al. [6] and have since been studied extensively in mathematics [8, 14]. The one-dimensional Fibonacci tiling is the canonical example of a cut-and-project quasicrystal [9]: generated by projecting a two-dimensional integer lattice onto the line $y=\varphi x$, it produces an aperiodic sequence of two tile types (L and S) with a precisely irrational frequency ratio $\varphi:1$.
Substitution systems and their combinatorial properties are surveyed in [12, 13]. The Fibonacci word as the canonical Sturmian sequence is treated in [3, 10, 11]. The Three-Distance (Steinhaus) theorem is due to Steinhaus [4] and Sós [5]. Pisot-Vijayaraghavan numbers and their role in substitution systems are discussed in [7, 15]. To our knowledge, Quasicryth is the first compressor to use a quasicrystalline substitution hierarchy as the primary parsing mechanism, and to prove that this hierarchy provides a structural compression advantage over periodic alternatives with bounded period.
2.4 Context Models and Hierarchy
The PAQ family [24] blends hundreds of adaptive bit-level context models, achieving state-of-the-art compression on structured data. Quasicryth takes a different approach: rather than many parallel context models, it uses a single deterministic quasicrystalline structure to expose a deep hierarchy of phrase-level contexts. The hierarchy context (3 bits per tile, free from the tiling) yields 8 specialised arithmetic coding sub-models per tile type at zero bitstream cost.
3. Method
Quasicryth processes input text through an eight-step pipeline. All design choices reflect the constraint that the tiling structure must be perfectly recoverable from a single phase parameter.
3.1 Step 1: Case Separation
The input byte stream is tokenised into words and punctuation tokens. Each token is classified as lowercase, titlecase, or uppercase, yielding a 3-symbol case flag stream encoded separately with a 18-bit-precision arithmetic coder (256-symbol adaptive model with periodic rescaling). The remaining pipeline operates on the fully lowercased token stream. On enwik9, case coding contributes 21,315,698 B ($\approx$2.13% of original), a necessary overhead for exact roundtrip.
3.2 Step 2: Word Tokenisation
The lowercased stream is split into word tokens: each token is either a maximal alphabetic run (with trailing whitespace absorbed) or a single non-alphabetic byte with trailing whitespace. This produces a sequence of $W$ word tokens. On enwik9, W=298,263,298W=298{,}263{,}298.
3.3 Step 3: Multi-Level Codebook Construction
Eleven frequency-ranked codebooks are built from the word token sequence, corresponding to phrase lengths $\{1,2,3,5,8,13,21,34,55,89,144\}$ — the first eleven Fibonacci numbers. Table 1 shows the codebook sizes. All n-gram codebooks are filtered to entries where every constituent word appears in the unigram codebook. For n-grams with $n\geq 13$, frequency counters are periodically pruned (singletons removed every 500 K entries) to prevent out-of-memory on large inputs. The serialised codebook is zlib-compressed; at enwik9 scale this occupies only 77,585 B.
Table 1: Codebook sizes (number of entries) by input size tier.
Level Phrase Input word count
len <<500K <<2M $\geq$10M
Unigram 1 4,000 8,000 8,000
Bigram 2 3,000 6,000 6,000
Trigram 3 1,000 2,000 2,000
5-gram 5 1,000 2,000 2,000
8-gram 8 500 1,000 1,000
13-gram 13 300 1,000 1,000
21-gram 21 100 500 500
34-gram 34 50 200 500
55-gram 55 0 0 200
89-gram 89 0 0 200
144-gram 144 0 0 100
Index encoding within the arithmetic coding stream uses a two-stage scheme: indices 0253253 are encoded directly (1 AC symbol); index $254$ signals an extended lookup (next symbol gives 254254508508); index $255$ signals an out-of-vocabulary escape.
3.4 Step 4: Quasicrystalline Word-Level Tiling
The one-dimensional Fibonacci quasicrystal is generated via the cut-and-project method. For tile position $k$ and phase $\theta\in[0,1)$:
$\text{tile}(k)=\begin{cases}L&\text{if }\lfloor(k+1+\theta)\,/\,\varphi\rfloor-\lfloor(k+\theta)\,/\,\varphi\rfloor=1\\ S&\text{otherwise}\end{cases}$ (1)
where $\varphi=(1+\sqrt{5})/2$ is the golden ratio. Each $L$ tile consumes 2 words (bigram lookup); each $S$ tile consumes 1 word (unigram lookup). The quasicrystalline matching rule — no two $S$ tiles adjacent — is enforced by merging any resulting $SS$ pair into an $L$ tile.
The compressor evaluates 32 equally spaced phases $\theta\in\{0/32,1/32,\ldots,31/32\}$, selecting the phase that maximises the scoring function:
$\text{score}=\sum_{k}b_{\ell(k)}$ (2)
where $\ell(k)$ is the deepest hierarchy level applicable at tile $k$ and $b_{\ell}$ are exponentially increasing bonuses: $b_{0}=3$ (unigram), $b_{1}=10$ (bigram), $b_{2}=20$ (trigram), $b_{3}=50$, $b_{4}=100$, $b_{5}=200$, $b_{6}=400$, $b_{7}=800$, $b_{8}=1{,}600$, $b_{9}=3{,}200$, $b_{10}=6{,}400$ (144-gram).
[…]
$S$: unigram (1 word)
$(L,S)\!\to\!\hat{L}_{1}$: trigram (3 w)
$L\!\to\!\hat{S}_{1}$: bigram (2 w)
$(\hat{L}_{1},\hat{S}_{1})\!\to\!\hat{L}_{2}$: 5-gram
$\hat{L}_{1}\!\to\!\hat{S}_{2}$: 3 words
Figure 1: Concrete example of Quasicryth tiling on the phrase “the quick brown fox jumps over the lazy” (8 words). Level 0: the Fibonacci cut-and-project rule assigns L tiles (2-word bigram positions, blue) and S tiles (1-word unigram positions, orange), yielding the sequence LSLLSL\;S\;L\;L\;S for these 8 words. Level 1: each $(L,S)$ pair is deflated into $\hat{L}_{1}$ (trigram position, teal); isolated $L$ tiles become $\hat{S}_{1}$. Level 2: the pair $(\hat{L}_{1},\hat{S}_{1})$ yields $\hat{L}_{2}$ (5-gram position, violet). Arrows trace the lookup cascade at word 0 (“the”): the encoder first attempts the 5-gram lookup (“the quick brown fox jumps”), then trigram (“the quick brown”), then bigram (“the quick”), falling back to unigram only if all longer lookups miss. Word 5 (“over”) reaches a trigram entry point but no 5-gram.
3.5 Step 5: Deep Substitution Hierarchy Construction
The inverse Fibonacci substitution (deflation) is applied iteratively to build a multi-scale hierarchy:
$\sigma^{-1}:(L,S)\to\text{super-}L,\quad L\text{ (isolated)}\to\text{super-}S$ (3)
At each deflation level $k$, the super-L tile at that level spans $F_{k+2}$ words, where $F_{n}$ is the $n$-th Fibonacci number. The function detect_deep_positions() traces each $L$ tile upward through the hierarchy: a tile at position $t_{i}$ can attempt a level-$k$ n-gram lookup if and only if it is the leftmost child of a super-L at every level up to $k$, and the span covers exactly $F_{k+2}$ words.
Figure 2: Deep hierarchy position detection. An $L$ tile at level 0 may qualify as the entry point of super-L supertiles at levels 1–9, enabling lookups in the trigram through 144-gram codebooks respectively. The qualifying condition is verified in a single upward pass through the pre-computed hierarchy array.
3.6 Step 6: Multi-Level Adaptive Arithmetic Coding
Encoding proceeds tile-by-tile. At each $L$ tile position, the deepest applicable hierarchy level is tried first. For level $k$ from $k_{\max}$ down to $1$: if the tile is at a super-L position at level $k$, encode a binary flag (hit/miss) via model lvl_flag[$k$]. On a hit, encode the codebook index via models lvl_idx[$k$] and lvl_ext[$k$], and advance the tile pointer by $F_{k+2}$ words. On a miss, fall through to the next lower level. After all deep levels fail, encode the bigram index via L_model[$c$] where $c\in\{0,\ldots,7\}$ is a 3-bit hierarchy context derived deterministically from the tile’s position in its parent, parent type, and grandparent type. $S$ tiles are encoded analogously via S_model[$c$].
Algorithm 1 summarises the encoding procedure. All arithmetic coding models use 256-symbol adaptive tables with 24-bit register precision and periodic frequency rescaling (halve all counts when the total exceeds 16,384).
Algorithm 1 Quasicryth encoding (payload stream)
0: Tile array $T$, word array $W$, codebooks $\mathcal{C}$
1: Build substitution hierarchy from $T$
2: Detect deep positions: $\text{can}[k][t_{i}]$ for $k=1\ldots k_{\max}$
[…]
4:while $i<|T|$ do
5:  if $T[i]=L$ then
6:   for $k=k_{\max}$ downto 11 do
7:    if $\text{can}[k][i]$ then
8:     Look up $W[i..i{+}F_{k+2}{-}1]$ in $\mathcal{C}_{k}$
9:     Encode flag $\mathbf{1}$ via lvl_flag[$k$]
10:     if hit then
11:      Encode index; $i\leftarrow i+F_{k+2}$; goto next
12:     end if
13:     Encode flag $\mathbf{0}$
14:    end if
15:   end for
16:   Encode bigram (W[i],W[i+1])(W[i],W[i{+}1]) via L_model[$c(i)$]
17:   $i\leftarrow i+2$
18:  else
19:   Encode unigram $W[i]$ via S_model[$c(i)$]
20:   $i\leftarrow i+1$
21:  end if
22:end while
3.7 Step 7: Separate bz2 Escape Stream
Out-of-vocabulary words (codebook misses) are collected into a separate buffer as length-prefixed raw byte sequences. The buffer is compressed with bzip2 (BWT + Huffman coding). This dual-stream architecture is critical: the arithmetic coding stream contains only compact integer indices and binary flags, while rare words compress efficiently under BWT due to their natural lexicographic clustering. On enwik9, the AC payload is 254,952,735 B and the bz2 escape stream is 83,537,358 B — combined they account for 93.6% of the total compressed size.
3.8 Step 8: Output Assembly
The compressed file format is:
Field Size
Magic (QT54) 4 B
Original size 4 B
Lowercased size 4 B
Word count 4 B
Phase (quantised $\theta$) 2 B
Flags 1 B
Case token count 4 B
Payload size 4 B
Case data size 4 B
Codebook size 4 B
Escape stream size 4 B
Payload variable
Case data variable
Codebook (zlib) variable
Escape stream (bz2) variable
MD5 checksum 16 B
The complete parsing structure — tile boundaries, hierarchy depths, deep n-gram positions, and model assignments — is implicit in the 2-byte phase value. The decompressor regenerates the Fibonacci tiling from $\theta$, rebuilds the substitution hierarchy, and decodes the AC stream identically, with no structural side-channel required.
4. Theoretical Analysis
This section develops the theoretical foundation of the paper. We state the main result first; the remainder of the section provides its proof, assembled from eight supporting theorems.
4.1 Main Result: Aperiodic Hierarchy Advantage
Theorem 1 (Aperiodic Hierarchy Advantage).
Let $T_{\text{fib}}$ be the Fibonacci quasicrystal tiling of $W$ words and $T_{\text{per}}$ any periodic tiling with period $p$ and collapse level $m^{*}=\lceil\log_{\varphi}p\rceil$. The Fibonacci tiling is the only infinite binary tiling satisfying all of the following simultaneously:
  1. (i)
    Non-collapse at every depth (Theorems 23): Both tile types are present at every hierarchy level $k\geq 0$, providing $F_{k}$-gram lookup positions for all $k$. Every periodic tiling collapses within $m^{*}=O(\log p)$ levels, yielding zero deep positions thereafter.
  2. (ii)
    Scale-invariant coverage (Theorem 9): The potential word coverage at each level satisfies $C(m)=P(m)\cdot F_{m}\to W\varphi/\sqrt{5}$, a constant independent of depth. Dictionary reuse capacity is maintained uniformly at all scales.
  3. (iii)
    Maximum codebook efficiency (Theorem 7, Corollary 8): Sturmian minimality (p(n)=n+1p(n)=n{+}1) guarantees exactly $F_{m}{+}1$ distinct tile-type patterns at level $m$, giving coverage efficiency $\eta_{m}=C_{m}/(F_{m}{+}1)$, maximal among aperiodic tilings.
  4. (iv)
    Bounded parsing overhead (Theorem 14): The per-word flag entropy satisfies $h_{\text{flags}}^{\text{fib}}\leq 1/\varphi\approx 0.618$ b/word, regardless of how many hierarchy levels are active. Access to all depths costs a convergent series of overhead.
  5. (v)
    Strict coding entropy advantage (Theorem 16): For any source with long-range phrase dependencies extending beyond level $m^{*}$ ($m$-LRPD, Definition 5):
    $H^{\text{fib}}(P)\;<\;H^{\text{per}}(P).$
As consequences: the net compression efficiency $\nu^{\text{fib}}(W)\to+\infty$ while $\nu^{\text{per}}$ is bounded (Corollary 15); coding redundancy decays super-exponentially $O(e^{-\varphi^{m}/\lambda})$ versus the fixed $\Omega(e^{-F_{m^{*}}/\lambda})$ of any periodic tiling (Theorem 18); and the per-entry dictionary value grows as $\Omega(\varphi^{m})$ with no periodic equivalent (Theorem 19).
Proof (assembly of supporting results).
Parts (i)–(v) are established in the following subsections. (i) from Theorems 2 and 34.44.5); (ii) from Theorem 94.12); (iii) from Theorem 74.10); (iv) from Theorem 144.15); (v) from Theorem 164.16). The ‘only’ claim follows from Corollary 8: any sequence with fewer than n+1n{+}1 distinct length-$n$ factors is periodic (collapses by (i)); any with strictly more is non-Sturmian (lower efficiency, violating (iii)). The three consequences follow from Corollary 15 and Theorems 1819. ∎
The proof infrastructure occupies the remainder of this section.
4.2 Substitution Matrix and Eigenvalue Analysis
The Fibonacci substitution $\sigma:L\to LS,\;S\to L$ has the substitution matrix
$M=\begin{pmatrix}1&1\\ 1&0\end{pmatrix}$ (4)
whose entry $M_{ij}$ counts how many tiles of type $i$ appear in $\sigma(j)$. The characteristic polynomial is $\lambda^{2}-\lambda-1=0$, with roots
$\lambda_{1}$ $=\varphi=\tfrac{1+\sqrt{5}}{2}\approx 1.618$ (5)
$\lambda_{2}$ $=\psi=\tfrac{1-\sqrt{5}}{2}\approx-0.618$ (6)
where $|\lambda_{2}|=1/\varphi<1$. By the Perron-Frobenius theorem [17], the dominant eigenvector of $M$ is
$\mathbf{v}_{1}=\frac{1}{\varphi+1}\begin{pmatrix}\varphi\\ 1\end{pmatrix}$ (7)
giving asymptotic tile frequencies $\text{freq}(L)=\varphi/(\varphi{+}1)\approx 0.618$ and $\text{freq}(S)=1/(\varphi{+}1)\approx 0.382$. The ratio $n_{L}/n_{S}=\varphi$ is irrational, which is the fundamental obstruction to periodicity.
4.3 The Inverse (Deflation) Substitution
Quasicryth applies the inverse Fibonacci substitution (desubstitution / deflation) at each hierarchy level. The inverse rule is:
$\sigma^{-1}:(L,S)\to\text{super-}L,\quad L\to\text{super-}S$ (8)
The corresponding matrix is
$M^{-1}=\begin{pmatrix}0&1\\ 1&-1\end{pmatrix}$ (9)
with eigenvalues $1/\varphi$ and $1/\psi$. Under $M^{-1}$, tile counts transform as
$n_{L}^{(k+1)}$ $=n_{S}^{(k)}$ (10)
$n_{S}^{(k+1)}$ $=n_{L}^{(k)}-n_{S}^{(k)}$ (11)
Starting from $N$ tiles at level 0, the number of supertiles at level $k$ is $N/\varphi^{k}$ (asymptotically), each spanning $F_{k+2}$ words of the original sequence.
Figure 3: Deep substitution hierarchy of the Fibonacci tiling. Each level is obtained by one application of the deflation rule $\sigma^{-1}$. Both $L$- and $S$-supertile types are present at every depth $k$, enabling $n$-gram codebook lookups at phrase lengths $F_{k+2}\in\{3,5,8,13,21,34,55,89,144,\ldots\}$ words.
4.4 Theorem: Fibonacci Hierarchy Never Collapses
Theorem 2 (Fibonacci Stability).
Let $T$ be the one-dimensional Fibonacci quasicrystal tiling generated by Eq. (1) for any phase $\theta$. For all $k\geq 0$, the level-$k$ supertile sequence contains both L-supertiles and S-supertiles.
Proof.
The frequency vector of the Fibonacci tiling is exactly the Perron-Frobenius eigenvector $\mathbf{v}_{1}$ (Eq. (7)). Under $(M^{-1})^{k}$:
$(M^{-1})^{k}\mathbf{v}_{1}=\left(\tfrac{1}{\varphi}\right)^{k}\mathbf{v}_{1}=\frac{1}{\varphi^{k}(\varphi{+}1)}\begin{pmatrix}\varphi\\ 1\end{pmatrix}$ (12)
The frequency vector is scaled by $1/\varphi^{k}$ but its direction is preserved exactly. The ratio $n_{L}^{(k)}/n_{S}^{(k)}=\varphi>0$ persists at every level $k$. Since $n_{L}^{(k)},n_{S}^{(k)}>0$ for all finite $k$, both supertile types are always present. ∎
Remark 1.
The stability expressed in Eq. (12) holds because $\mathbf{v}_{1}$ is the only eigenvector of $M^{-1}$ with eigenvalue $|\lambda|<1$. The second eigenvalue $1/\psi=\varphi\cdot\text{sign}(\psi)$ satisfies $|1/\psi|=\varphi>1$, so any perturbation away from $\mathbf{v}_{1}$ grows exponentially under deflation. The Fibonacci tiling is precisely the fixed point that avoids this divergence, because $\varphi$ is a Pisot-Vijayaraghavan number (see Section 4.6).
4.5 Theorem: All Periodic Tilings Collapse
Theorem 3 (Periodic Collapse).
Let $T$ be a periodic tiling with period $p$ and rational $L$-frequency $r=n_{L}/p$. Then there exists a finite level $k^{*}\leq\lceil\log_{\varphi}p\rceil$ such that the level-$k^{*}$ supertile sequence contains only one tile type.
Proof.
Decompose the frequency vector in the eigenbasis of $M$:
$\mathbf{v}=\alpha\mathbf{v}_{1}+\beta\mathbf{v}_{2}$ (13)
where $\mathbf{v}_{2}$ is the eigenvector for $\lambda_{2}=\psi$. Since $\mathbf{v}$ has rational entries and $\mathbf{v}_{1}$ has irrational entries (containing $\sqrt{5}$), we must have $\beta\neq 0$. Under $(M^{-1})^{k}$:
$(M^{-1})^{k}\mathbf{v}=\alpha\varphi^{-k}\mathbf{v}_{1}+\beta\psi^{-k}\mathbf{v}_{2}$ (14)
Since $|\psi^{-1}|=\varphi>1$, the second component grows as $\varphi^{k}$, dominating for large $k$. The frequency trajectory $n_{L}^{(k)}/n_{S}^{(k)}$ therefore diverges from $\varphi$: it oscillates and eventually sends one count to zero (or negative, indicating collapse has already occurred). The collapse timescale is determined by when $\beta\varphi^{k}|\psi^{-k}|$ exceeds the first component:
$k^{*}\approx\frac{\log p}{\log\varphi}$ (15)
since the deviation $|\beta|\propto 1/p$ for the best periodic approximant of $\varphi$ with period $p$. ∎
Corollary 4.
For Period-5 ($p=5$), collapse occurs at level $k^{*}\approx\log 5/\log\varphi\approx 3.3$, confirmed experimentally at level 4 (Section 4.7).
4.6 Pisot-Vijayaraghavan Property
Definition 1 (Pisot-Vijayaraghavan number [7]).
A real algebraic integer $\beta>1$ is a Pisot-Vijayaraghavan (PV) number if all its algebraic conjugates have absolute value strictly less than 1.
The golden ratio $\varphi=(1+\sqrt{5})/2$ is a PV number: its conjugate $\psi=(1-\sqrt{5})/2$ satisfies $|\psi|=0.618<1$. This is the minimal PV number (degree 2).
The PV property has two critical consequences for Quasicryth:
  1. 1.
    Recognisability. The Fibonacci substitution is recognisable [16]: every bi-infinite Fibonacci word has a unique desubstitution. This is precisely the property that allows Quasicryth’ decompressor to reconstruct the tiling hierarchy unambiguously from the phase alone.
  2. 2.
    Self-similarity. The fixed point of $\sigma$ (the infinite Fibonacci word) is also the fixed point of $\sigma^{k}$ for all $k>0$. This means the supertile sequences at all levels are simply scaled copies of the original tiling — the compressor is literally operating on the same structure at every hierarchical scale.
4.7 Worked Example: Period-5 Collapse
Period-5 = LLSLS repeating (3L, 2S per period, $r=0.600$) is the Fibonacci approximant of order 5: the continued fraction $\varphi=[1;1,1,1,\ldots]$ truncated at the 5th convergent gives $p_{4}/q_{4}=3/5=0.600$.
Level Supertiles $n_{L}$ $n_{S}$
0 LLSLS (period) 3 2
1 SLL 2 1
2 SSL 1 2
3 L 1 0
4 S 0 1
The $L$-frequency trajectory $0.600\to 0.667\to 0.333\to 1.0\to 0.0$ oscillates with growing amplitude, driven by the $\varphi^{k}$ factor in the second eigenvector component. At level 4, all tiles are $S$: there are zero super-L positions, hence zero positions available for 13-gram, 21-gram, 34-gram, 55-gram, 89-gram, or 144-gram lookups. This collapse is not a numerical artifact but an exact algebraic consequence of the rational frequency $3/5\neq\varphi$.
4.8 Connection to Weyl’s Equidistribution Theorem
Theorem 5 (Weyl [2]).
For any irrational $\alpha$ and any $a,b)\subset[0,1)$,
$\lim_{N\to\infty}\frac{1}{N}\sum_{k=0}^{N-1}\mathbf{1}_{\{k\alpha\bmod 1\in[a,b)\}}=b-a$ (16)
The Quasicryth tiling (Eq. (1)) is equivalent to the irrational rotation by angle $1/\varphi$ on the unit circle: assign $L$ to position $k$ if and only if $k/\varphi+\theta\pmod{1}\in[0,1/\varphi)$. By Weyl’s theorem with $\alpha=1/\varphi$ (irrational), the $L$-frequency converges to exactly $1/\varphi$ at every scale.
Crucially, the supertile sequence at level $k$ is again an irrational rotation — with angle $1/\varphi^{k+1}$, also irrational. Equidistribution therefore holds at every hierarchy level, guaranteeing uniform tile density throughout the hierarchy. Periodic tilings correspond to rational $\alpha=p/q$, which visit only $q$ distinct points on the circle; equidistribution fails at scales larger than the period.
4.9 The Three-Distance Theorem and Sturmian Structure
Theorem 6 (Three-Distance Theorem [4, 5]).
For any irrational $\alpha$ and any $N\geq 1$, the $N$ points $\{k\alpha\bmod 1:0\leq k<N\}$ partition the circle $[0,1)$ into gaps of at most three distinct lengths, the largest being the sum of the other two.
For $\alpha=1/\varphi$, the three gap lengths correspond to the three local configurations in the Fibonacci tiling: $LS$ (long gap), an isolated $L$ (medium gap), and $S$ (short gap). This structural richness is preserved at every hierarchy level, and it is precisely what differentiates the Fibonacci tiling from its periodic approximants, which have only two gap lengths.
The Fibonacci word is the canonical Sturmian sequence. Following Morse and Hedlund [3]:
Definition 2 (Sturmian sequence).
An infinite binary sequence $u$ is Sturmian if it is (1) aperiodic and (2) balanced: for any two factors $w$, $w^{\prime}$ of equal length, $|\#_{L}(w)-\#_{L}(w^{\prime})|\leq 1$.
A sequence is Sturmian if and only if it is an irrational rotation coding [10]. The Sturmian property has three direct implications for compression:
  1. 1.
    Balance. Any two windows of equal length have L-counts differing by at most 1. This means every n-gram codebook entry is equally useful at any position in the input: there are no “dense” or “sparse” regions for long-range lookups.
  2. 2.
    Aperiodicity. The sequence is never eventually periodic, so the deflation hierarchy never produces an all-S or all-L level (Theorem 2). Codebook positions exist at all depths indefinitely.
  3. 3.
    Minimal factor complexity. A Sturmian sequence of alphabet size 2 has exactly $n+1$ distinct factors of length $n$, the minimum possible for an aperiodic sequence [3]. For the Fibonacci tiling, this means the $n$-gram codebooks have coverage efficiency that is maximal in this setting: with $n+1$ distinct $n$-tuples of tile types, $n$-gram codebook entries match at the maximum fraction of eligible positions.
No periodic sequence satisfies all three conditions: any periodic sequence is eventually periodic (violating condition 2) and has bounded factor complexity (violating condition 3 for large $n$).
4.10 Theorem: Sturmian Codebook Efficiency
The minimal factor complexity established above has a precise quantitative implication for codebook coverage that we now formalise.
Theorem 7 (Sturmian Codebook Efficiency).
Let $u$ be the Fibonacci word over $\{L,S\}$. For every $n\geq 1$, the number of distinct length-$n$ factors of $u$ is exactly $n+1$, the minimum possible for any aperiodic binary sequence (Morse–Hedlund theorem [3]). Consequently, at hierarchy level $m$ a codebook of $C_{m}$ entries covers the fraction
$\eta_{m}\;=\;\frac{\min(C_{m},\;F_{m}+1)}{F_{m}+1}$ (17)
of all distinct $F_{m}$-gram tile-type patterns. For any aperiodic non-Sturmian binary tiling, $p(F_{m})\geq F_{m}+2$, giving strictly lower efficiency $\eta_{m}\leq C_{m}/(F_{m}+2)<C_{m}/(F_{m}+1)$ for the same codebook size.
Proof.
The Morse–Hedlund theorem [3] states that an infinite binary sequence is aperiodic if and only if $p(n)\geq n+1$ for all $n\geq 1$, and Sturmian if and only if equality holds throughout. The Fibonacci word is the canonical Sturmian sequence over $\{L,S\}$ (see also [11]), so $p(n)=n+1$ for all $n$. The supertile sequence at each hierarchy level $k$ is the image of the original Fibonacci word under $(\sigma^{-1})^{k}$; since $\sigma^{-1}$ preserves the PV fixed-point structure (Section 4.6), every level is also Sturmian and satisfies $p(n)=n+1$. A codebook of $C_{m}$ entries covers $\min(C_{m},F_{m}+1)$ of the $F_{m}+1$ distinct $F_{m}$-letter tile words, giving (17). Any aperiodic non-Sturmian sequence has $p(F_{m})\geq F_{m}+2$, reducing coverage to at most $C_{m}/(F_{m}+2)$. ∎
Corollary 8 (Fibonacci as Goldilocks Tiling).
Among all infinite aperiodic binary sequences, the Fibonacci word simultaneously achieves:
  1. 1.
    Hierarchy non-collapse at every depth (Theorem 2): both tile types persist at every level, providing $n$-gram lookup positions indefinitely.
  2. 2.
    Maximum codebook efficiency at every level (Theorem 7): only $F_{m}+1$ distinct tile-type patterns exist at level $m$, so a codebook of $C_{m}$ entries achieves hit-rate density $C_{m}/(F_{m}+1)$, maximal in this setting.
Any sequence with factor complexity <n+1<n+1 is periodic and collapses (Theorem 3). Any sequence with complexity >n+1>n+1 is non-Sturmian and incurs strictly lower codebook efficiency. The Fibonacci tiling is therefore characterised as the only aperiodic sequence satisfying both properties at every scale simultaneously.
The practical consequence is that the $F_{m}$-gram codebook fills with genuinely distinct phrase patterns — there are only $F_{m}+1$ distinct tile-type contexts, so a codebook of that size covers the full structural diversity of the level. This is why empirical hit rates (Table 7) do not decay to zero as $m$ grows: Sturmian minimality prevents the structural variety from exploding exponentially with level depth.
4.11 Information-Theoretic Bound on the Aperiodic Advantage
At hierarchy level $k$, the Fibonacci tiling provides approximately
$P_{\text{fib}}(k)\approx\frac{W}{\varphi^{k+1}(\varphi+1)}$ (18)
super-L positions for $F_{k+2}$-gram lookup, where $W$ is the total word count. Any periodic tiling with collapse level $k^{*}$ provides $P_{\text{per}}(k)=0$ for all $k\geq k^{*}$.
Each $n$-gram hit at level $k$ saves approximately
$\Delta(k)=F_{k+2}\cdot h-\log_{2}|\mathcal{C}_{k}|\quad\text{bits}$ (19)
where $h$ is the per-word entropy of the arithmetic coder and $|\mathcal{C}_{k}|$ is the codebook size at level $k$. The total aperiodic advantage over a periodic tiling with collapse at $k^{*}$ is
$\text{Adv}=\sum_{k=k^{*}}^{k_{\max}}P_{\text{fib}}(k)\cdot r(k)\cdot\Delta(k)$ (20)
where $r(k)$ is the hit rate at level $k$. For fixed $k_{\max}$, this scales as $O(W)$ (linear in word count).
The superlinear jump observed empirically arises when a new hierarchy level $k_{\max}+1$ activates: at that point, $P_{\text{fib}}(k_{\max}+1)\approx W/\varphi^{k_{\max}+2}(\varphi+1)$ new positions become available, each contributing $\Delta(k_{\max}+1)$ savings. The total advantage then grows as $O(W/\varphi^{k_{\max}+2})$ for the new level, on top of the already linear contributions from levels $k^{*}\ldots k_{\max}$.
At enwik9 (W=298,263,298W=298{,}263{,}298 words), levels 8 and 9 activate for the first time, providing:
$P_{\text{fib}}(8)$ $\approx\frac{298M}{\varphi^{9}(\varphi{+}1)}\approx 2{,}400$ (21)
$P_{\text{fib}}(9)$ $\approx\frac{298M}{\varphi^{10}(\varphi{+}1)}\approx 1{,}480$ (22)
[…]
4.12 Theorem: Golden Compensation
The following theorem establishes a scale-invariant coverage property of the Fibonacci hierarchy.
Definition 3 (Fibonacci n-gram level).
For a Fibonacci number $F_{m}$ ($m\geq 3$), the $F_{m}$-gram level consists of all super-L tile positions whose phrase spans exactly $F_{m}$ consecutive words. The number of such positions in a Fibonacci tiling of $W$ words is
$P(m)\;=\;\frac{W}{\varphi^{m-1}}$ (23)
asymptotically, where the index $m$ runs through the Fibonacci indices $m=4,5,6,7,8,9,10,11,12$ corresponding to phrase lengths $F_{m}\in\{3,5,8,13,21,34,55,89,144\}$ words.
Theorem 9 (Golden Compensation).
In a Fibonacci quasicrystal tiling of $W$ words, the potential word coverage at the $F_{m}$-gram level — i.e., the total words that could be encoded by super-L lookups at that level — is
$\begin{split}C(m)&\;:=\;P(m)\cdot F_{m}\;=\;\frac{W\cdot F_{m}}{\varphi^{m-1}}\\ &\;\longrightarrow\;\frac{W\varphi}{\sqrt{5}}\quad\text{as }m\to\infty.\end{split}$ (24)
The limit $W\varphi/\sqrt{5}\approx 0.724\,W$ is independent of the hierarchy level $m$.
Proof.
By Binet’s formula, $F_{m}=(\varphi^{m}-\psi^{m})/\sqrt{5}$, where $|\psi|=1/\varphi<1$. Then
$\frac{F_{m}}{\varphi^{m-1}}=\frac{\varphi^{m}-\psi^{m}}{\sqrt{5}\,\varphi^{m-1}}=\frac{\varphi-(\psi/\varphi)^{m}\,\psi}{\sqrt{5}}\;\longrightarrow\;\frac{\varphi}{\sqrt{5}},$
since $(\psi/\varphi)^{m}\to 0$. For all finite $m$ the exact value is $C(m)=W(\varphi^{m}-\psi^{m})/(\sqrt{5}\,\varphi^{m-1})$. ∎
The theorem rests on a cancellation between two competing effects:
  • Position-count decay: deeper levels have fewer super-L positions, with $P(m)\propto\varphi^{-(m-1)}$.
  • Phrase-length growth: deeper levels encode longer phrases, with $F_{m}\propto\varphi^{m}$.
The two $\varphi$-exponentials cancel exactly, because $F_{m}\sim\varphi^{m}/\sqrt{5}$ (Binet) and $\varphi^{2}=\varphi+1$ (defining identity of the golden ratio). This cancellation is a structural consequence of $\varphi$ being the Perron-Frobenius eigenvalue of the Fibonacci substitution matrix.
Corollary 10 (Word-Weighted Contribution).
The word-weighted compression contribution of the $F_{m}$-gram level is
$P(m)\cdot r(m)\cdot F_{m}\;\approx\;r(m)\cdot\frac{W\varphi}{\sqrt{5}},$ (25)
where $r(m)\in[0,1]$ is the codebook hit rate at level $m$. All structural variation across hierarchy levels is therefore encoded entirely in $r(m)$; the positional and phrase-length factors are architecture-neutral.
Corollary 11 (Periodic Tilings Contribute Zero).
For any periodic tiling with collapse level $m^{*}$, $P(m)=0$ for all $m\geq m^{*}$, so the contribution $C(m)=0$ for all deep levels regardless of how large $W$ grows. The Fibonacci hierarchy, by contrast, maintains $C(m)\to W\varphi/\sqrt{5}$ at every level.
Empirically, Table 7 confirms the near-constant potential coverage: at enwik9 (W=298,263,298W=298{,}263{,}298), the theoretical $C(m)=W\varphi/\sqrt{5}\approx 215{,}840{,}000$ words per level is modulated by hit rates $r(m)$ decreasing from 3.9% at 13-gram to 0.06% at 144-gram, giving the observed word-weighted shares of 65.7% down to 1.1% (Fig. 5).
4.13 Theorem: Level Activation Threshold
Theorem 12 (Activation Threshold).
The $F_{m}$-gram codebook first produces at least one useful entry when
$W\;\geq\;W^{*}_{m}\;:=\;\frac{T_{m}\cdot\varphi^{m-1}}{r_{m}},$ (26)
where $T_{m}$ is the minimum number of codebook entries required for the level to contribute ($T_{m}\geq 1$), and $r_{m}$ is the expected hit rate.
Proof.
The expected number of $F_{m}$-gram hits is $P(m)\cdot r_{m}=W\,r_{m}/\varphi^{m-1}$. Setting this $\geq T_{m}$ and solving for $W$ gives (26). ∎
Table 2: Predicted activation thresholds $W^{*}_{m}$ for the two deepest levels (89-gram and 144-gram), using empirical hit rates from enwik9 and $T_{m}=1$.
Level $F_{m}$ $\varphi^{m-1}$ Hit rate $r_{m}$ $W^{*}_{m}$
89-gram $F_{11}$ $\approx 123$ 0.10% $\approx 123{,}000$ words
144-gram $F_{12}$ $\approx 199$ 0.06% $\approx 332{,}000$ words
The threshold for at least one hit is easily exceeded; the more practically relevant threshold is when the codebook accumulates enough repeated $F_{m}$-grams to pass pruning and yield non-trivial entries. Empirically, the 89-gram and 144-gram codebooks are empty at enwik8 (W=27,700, [^0] W=27{,}700{,}000, 0 hits) and active at enwik9 (W=298,263,298W=298{,}263{,}298, giving 2,396 and 945 hits respectively), consistent with the codebook activation scaling as $O(\varphi^{m-1})$.
4.14 Theorem: Piecewise-Linear Aperiodic Advantage
Combining Theorems 9 and 12, we characterise the growth of the aperiodic advantage $A(W)$ (Fibonacci vs. best periodic approximant) as a function of corpus size.
Theorem 13 (Piecewise-Linear Advantage).
Let $m^{*}$ be the collapse level of the periodic baseline, and let $m_{1}<m_{2}<\cdots$ be the sequence of Fibonacci indices for which $W>W^{*}_{m_{i}}$ (active deep levels). Then
$A(W)\;=\;\sum_{m=m^{*}}^{m_{\max}(W)}\frac{W}{\varphi^{m-1}}\,r(m)\,\Delta(m),$ (27)
where $\Delta(m)=F_{m}\cdot h-\log_{2}|\mathcal{C}_{m}|$ is the per-hit saving and $m_{\max}(W)=\max\{m:W\geq W^{*}_{m}\}$.
Between consecutive activation thresholds $W^{*}_{m}$ and $W^{*}_{m+1}$, $A(W)$ is strictly linear in $W$ with slope
$s_{m}\;=\;\sum_{m^{\prime}=m^{*}}^{m}\frac{r(m^{\prime})\,\Delta(m^{\prime})}{\varphi^{m^{\prime}-1}}.$ (28)
At each threshold $W=W^{*}_{m+1}$, the slope increases by $r(m+1)\,\Delta(m+1)/\varphi^{m}>0$. Hence $A(W)$ is a piecewise-linear, convex function of $W$ with monotonically increasing slopes.
The empirically measured advantage values confirm this structure (Table 3).
Table 3: Aperiodic advantage (Fibonacci vs. Period-5) at each scale, with the active deep levels and whether a slope increase occurs at that scale, as predicted by Theorem 13.
Scale Advantage Active deep levels Slope change
3 MB 1,372 B 13g, 21g, 34g
10 MB 5,807 B +55g activated $s_{55g}$
100 MB 40,385 B (same) none
1 GB 1,349,371 B +89g, +144g activated $s_{89g}+s_{144g}$
[…]
4.15 Theorem: Aperiodic Parsing Dominance
We now bound the per-word cost of accessing all hierarchy depths.
Definition 4 (Per-word flag entropy).
The per-word flag entropy of a tiling up to depth $k_{\max}$ is
$h_{\text{flags}}(k_{\max})\;:=\;\frac{1}{W}\sum_{k=1}^{k_{\max}}P(k)\cdot H_{b}(r_{k}),$ (29)
where $H_{b}(r)=-r\log_{2}r-(1-r)\log_{2}(1-r)$ is the binary entropy function.
Theorem 14 (Convergent Flag Overhead).
For the Fibonacci quasicrystal tiling with $P(k)=W/\varphi^{k+1}(\varphi+1)$:
$h_{\text{flags}}^{\text{fib}}\;=\;\lim_{k_{\max}\to\infty}h_{\text{flags}}(k_{\max})\;\leq\;\frac{1}{\varphi}.$ (30)
Proof.
Since $H_{b}(r)\leq 1$ for all $r\in[0,1]$:
$h_{\text{flags}}^{\text{fib}}$ $\;\leq\;\frac{1}{\varphi+1}\sum_{k=1}^{\infty}\varphi^{-k}\;=\;\frac{1}{\varphi+1}\cdot\frac{1}{\varphi-1}$
$\;=\;\frac{1}{\varphi^{2}-1}\;=\;\frac{1}{\varphi},$
using $\varphi^{2}=\varphi+1$ in the final step. The series converges because position counts decay geometrically with ratio $1/\varphi<1$. ∎
Each additional hierarchy level contributes flag overhead $\varphi$ times smaller than the previous, since the position counts decay by the same factor.
Corollary 15 ($H_{\text{aperiodic}}<H_{\text{periodic}}$ in efficiency).
Define the per-word net efficiency
$\nu(W)$ $\;:=\;h_{\text{saved}}(W)-h_{\text{flags}}(W),$
$h_{\text{saved}}$ $:=\frac{1}{W}\sum_{k}P(k)\,r_{k}\,\Delta_{k}.$ (31)
Then for any periodic tiling with collapse level $k^{*}$, there exists $W_{0}$ such that for all $W\geq W_{0}$:
$\nu^{\text{fib}}(W)\;>\;\nu^{\text{per}},\qquad\lim_{W\to\infty}\nu^{\text{fib}}(W)=+\infty.$ (32)
Proof.
By Theorem 13, $h_{\text{saved}}^{\text{fib}}(W)$ grows without bound as successive levels activate, each adding a strictly positive $O(W)$ term. By Theorem 14, $h_{\text{flags}}^{\text{fib}}$ is bounded above by $1/\varphi$. Therefore $\nu^{\text{fib}}(W)=h_{\text{saved}}^{\text{fib}}(W)-h_{\text{flags}}^{\text{fib}}\to+\infty$. For any periodic tiling: collapse at $k^{*}$ means $h_{\text{saved}}^{\text{per}}$ is fixed for all $W$ (no new levels activate), so $\nu^{\text{per}}$ is a bounded constant. Equation (32) follows. ∎
The contrast is stark: the Fibonacci tiling pays at most $1/\varphi$ bits per word in flag overhead, regardless of how many levels are active, while the compression benefit grows without bound. For any periodic tiling, the overhead is smaller (fewer flag positions at deep levels) but the benefit is also smaller — permanently capped at the collapse level. The two conditions $\nu^{\text{fib}}\to+\infty$ and $\nu^{\text{per}}$ bounded constitute a formal sense in which the overhead structure of aperiodic tilings is provably superior under the stated conditions.
4.16 Theorem: Strict Coding Entropy for Long-Range Sources
For sources with genuine long-range phrase dependencies, the per-word coding entropy is lower under the Fibonacci hierarchy than under any periodic alternative.
Definition 5 (Long-range phrase dependency).
A stationary ergodic source $P$ over word sequences is $m$-long-range phrase dependent ($m$-LRPD) if
$h_{F_{m}}(P)\;<\;h_{F_{m-1}}(P),$ (33)
where $h_{k}(P):=H(X_{0}\mid X_{-1},\ldots,X_{-k+1})$ is the entropy rate under optimal $k$-gram prediction. Equivalently, the $F_{m}$-word context provides strictly more information about the current word than the $F_{m-1}$-word context.
Theorem 16 (Strict Coding Entropy Inequality).
Let $P$ be an $m$-LRPD source for some $m>m^{*}$, where $m^{*}$ is the collapse level of a periodic tiling. In the limit of sufficient codebook capacity,
$H^{\text{fib}}(P)\;\leq\;h_{F_{m}}(P)\;<\;h_{F_{m^{*}}}(P)\;\leq\;H^{\text{per}}(P).$ (34)
In particular, $H^{\text{fib}}(P)<H^{\text{per}}(P)$ strictly.
Proof.
Conditioning on additional context never increases entropy: for $k<k^{\prime}$,
$h_{k}(P)$ $=H(X_{0}\mid X_{-1},\ldots,X_{-k+1})$
$\;\geq\;H(X_{0}\mid X_{-1},\ldots,X_{-k^{\prime}})=h_{k^{\prime}}(P),$ (35)
with strict inequality iff $X_{-k:-(k^{\prime}-1)}$ carries additional information about $X_{0}$; so $h_{k}$ is non-increasing in $k$.
Left inequality.  The Fibonacci tiling at level $m$ provides super-L positions with $F_{m}$-gram spans. A codebook approximating $P(X_{0}\mid X_{-1:-(F_{m}-1)})$ drives the per-word coding cost at those positions to $h_{F_{m}}(P)$, so $H^{\text{fib}}(P)\leq h_{F_{m}}(P)$.
Strict central inequality.  By $m$-LRPD: $h_{F_{m}}(P)<h_{F_{m-1}}(P)$. Since $m>m^{*}$, we have $F_{m-1}\geq F_{m^{*}}$, so $h_{F_{m-1}}(P)\leq h_{F_{m^{*}}}(P)$ by (35). Combining: $h_{F_{m}}(P)<h_{F_{m^{*}}}(P)$ strictly.
Right inequality.  The periodic tiling collapses at $m^{*}$: no super-L positions exist beyond depth $m^{*}$, so the coder is limited to contexts of length $\leq F_{m^{*}}$, giving $H^{\text{per}}(P)\geq h_{F_{m^{*}}}(P)$. ∎
Corollary 17 (Natural language is $m$-LRPD).
The non-zero hit counts at levels 13g–144g in Table 7 (e.g. 652,124 hits at enwik9) demonstrate directly that Wikipedia text is $m$-LRPD for $m$ corresponding to phrase lengths up to at least 144 words. For any periodic tiling with collapse at $m^{*}\leq 4$ (period $p\leq\varphi^{4}\approx 7$, phrase length $\leq 5$ words), the strict inequality $H^{\text{fib}}(P)<H^{\text{per}}(P)$ holds on Wikipedia-class corpora. The 1,349,371 B advantage at enwik9 is the byte-level expression of the entropy gap $h_{F_{m^{*}}}(P)-h_{F_{m}}(P)$ summed over 298 M word positions.
The theorem separates two distinct sources of the aperiodic advantage:
  • Structural: the Fibonacci hierarchy provides more positions at every deep level (Theorems 2 and 9).
  • Information-theoretic: for $m$-LRPD sources, those positions operate at strictly lower entropy per word, so even an equal number of positions would yield strictly more compression.
For sources with no long-range phrase structure ($h_{k}=\mathrm{const}$ for all $k\geq k_{0}\leq F_{m^{*}}$), the two tilings achieve equal coding entropy and the advantage reduces entirely to the combinatorial terms of Theorem 13.
4.17 Compression Bounds: Redundancy and Dictionary Efficiency
We sharpen the previous results with two quantitative bounds: a redundancy bound showing super-exponential decay in hierarchy depth, and a dictionary efficiency bound showing that each new Fibonacci level contributes codebook entries with exponentially growing per-entry compression value.
Redundancy Bound
Definition 6 (Exponentially mixing source).
A stationary ergodic source $P$ is exponentially mixing with parameters $(C,\lambda)$ if
$I(X_{0};\,X_{-k})\;\leq\;C\,e^{-k/\lambda},\qquad k\geq 1.$ (36)
Natural language exhibits approximately exponential mixing: long-range mutual information $I(X_{0};X_{-k})$ decays rapidly in $k$ [1].
Theorem 18 (Fibonacci Redundancy Bound).
Let $P$ be exponentially mixing with parameters $(C,\lambda)$. The $k$-gram coding redundancy $R_{k}(P):=h_{k}(P)-h(P)$ satisfies
$R_{k}(P)\;\leq\;\frac{C\,e^{-k/\lambda}}{1-e^{-1/\lambda}}.$ (37)
For the Fibonacci hierarchy at level $m$ ($k=F_{m}$):
$R_{m}^{\text{fib}}(P)\;=\;O\!\left(e^{-F_{m}/\lambda}\right)\;=\;O\!\left(e^{-\varphi^{m}/(\lambda\sqrt{5})}\right),$ (38)
decaying super-exponentially in $m$ (exponential in $\varphi^{m}$). For any periodic tiling with collapse at $m^{*}$:
$R^{\text{per}}(P)\;=\;R_{F_{m^{*}}}(P)\;=\;\Omega\!\left(e^{-F_{m^{*}}/\lambda}\right)>0.$ (39)
The redundancy ratio satisfies
$\frac{R_{m}^{\text{fib}}(P)}{R^{\text{per}}(P)}\;=\;O\!\left(e^{-(F_{m}-F_{m^{*}})/\lambda}\right)\;\xrightarrow{m\to\infty}\;0.$ (40)
Proof.
Since $h_{k}(P)$ is non-increasing and $h_{\infty}(P)=h(P)$:
$R_{k}(P)$ $=\sum_{n=k}^{\infty}\bigl[h_{n}(P)-h_{n+1}(P)\bigr]$
$=\sum_{n=k}^{\infty}I\!\left(X_{0};\,X_{-n}\mid X_{-1},\ldots,X_{-n+1}\right)$
$\leq\sum_{n=k}^{\infty}I(X_{0};\,X_{-n})$
$\leq C\sum_{n=k}^{\infty}e^{-n/\lambda}\;=\;\frac{C\,e^{-k/\lambda}}{1-e^{-1/\lambda}},$
using the data processing inequality in the second step and (36) in the third. Setting $k=F_{m}$ and using $F_{m}=\varphi^{m}/\sqrt{5}+O(|\psi|^{m})$ gives (38). The lower bound (39) holds for any $m$-LRPD source at scale $k=F_{m^{*}}$, by Definition 5. ∎
The super-exponential decay $O(e^{-\varphi^{m}/(\lambda\sqrt{5})})$ comes from the doubly-exponential growth of $F_{m}$: each Fibonacci level removes an exponentially larger slice of residual redundancy than the previous level. The periodic tiling is permanently locked at residual redundancy $R_{F_{m^{*}}}(P)$, while the Fibonacci hierarchy drives redundancy to zero in the limit.
Dictionary Efficiency
Theorem 19 (Exponential Dictionary Efficiency).
Define the per-entry compression gain at level $m$ as the expected bits saved per codebook hit:
$E_{m}\;:=\;F_{m}\,\bar{h}\;-\;\log_{2}C_{m},$ (41)
where $\bar{h}$ is the per-word arithmetic coding entropy under the baseline (unigram) model. Since codebook sizes $C_{m}$ are bounded (Table 1: $C_{m}\leq 8{,}000$ for all $m$, decreasing to 100 at the 144-gram level) and $F_{m}\sim\varphi^{m}/\sqrt{5}$ grows exponentially:
$E_{m}\;\sim\;F_{m}\,\bar{h}\;=\;\Omega(\varphi^{m}).$ (42)
The total dictionary value
$\mathcal{V}(k_{\max})\;:=\;\sum_{m=1}^{k_{\max}}C_{m}\,E_{m}\;=\;\Omega(\varphi^{k_{\max}})$ (43)
grows exponentially in $k_{\max}$ for the Fibonacci hierarchy, whereas for any periodic tiling with collapse at $m^{*}$: $\mathcal{V}^{\text{per}}=O(\varphi^{m^{*}})$ is a fixed constant. The ratio $\mathcal{V}^{\text{fib}}(k_{\max})/\mathcal{V}^{\text{per}}=\Omega(\varphi^{k_{\max}-m^{*}})\to\infty$.
Proof.
Since $C_{m}\leq 8{,}000$ for all $m$, we have $\log_{2}C_{m}\leq 13$. For all $m$ such that $F_{m}\,\bar{h}>13$ (i.e., $m\geq m_{0}$ for some fixed $m_{0}$ depending only on $\bar{h}$), $E_{m}>0$ and $E_{m}\geq F_{m}\,\bar{h}-13\sim F_{m}\,\bar{h}$. Then
$\mathcal{V}(k_{\max})\geq C_{k_{\max}}E_{k_{\max}}=\Omega(C_{k_{\max}}\varphi^{k_{\max}})=\Omega(\varphi^{k_{\max}}),$
since $C_{k_{\max}}\geq 1$. The periodic bound $\mathcal{V}^{\text{per}}=O(\varphi^{m^{*}})$ follows from terminating the sum at $m^{*}$ and bounding $F_{m}\leq F_{m^{*}}$ for $m\leq m^{*}$. ∎
Concretely: the 144-gram level ($m=9$, $F_{9}=144$, $C_{9}=100$) has per-entry gain $E_{9}=144\bar{h}-\log_{2}(100)\approx 144\bar{h}-6.6$ bits, encoding 144 words in place of a single AC symbol. This is $\approx 72\times$ the per-entry gain of the bigram level ($E_{1}\approx 2\bar{h}$), as expected from $F_{9}/F_{1}=72$. Each new Fibonacci level doubles the per-entry value in this sense.
4.18 Summary of Mathematical Properties
Table 4 contrasts the key mathematical properties of Fibonacci quasicrystal tilings versus all periodic tilings.
Table 4: Mathematical properties of Fibonacci quasicrystal versus periodic tilings, and their implications for compression.
Property Fibonacci Periodic ($p$)
Eigenvalue of $M$ $\varphi$ (PV, irrat.) rational freq.
$n_{L}/n_{S}$ at level $k$ $\varphi$ (constant) degenerates
Hierarchy depth $\infty$ $O(\log p)$
Factor complexity n+1n{+}1 (Sturmian) $\leq p$ (periodic)
Codebook eff. $\eta_{m}$ $C_{m}/(F_{m}{+}1)$ (max) 0 for $m\geq m^{*}$
Equidistribution Weyl (all scales) fails at $\geq p$
Recognisability unique deflation ambiguous/degen.
Deep n-gram pos. $O(W/\varphi^{k})$, all $k$ 0 for $k\geq k^{*}$
Coverage $P(m)\cdot F_{m}$ $\to W\varphi/\sqrt{5}$ (const.) 0 for $m\geq m^{*}$
Advantage $A(W)$ piecewise-linear, convex baseline (0)
Flag overhead $h_{\text{flags}}$ $\leq 1/\varphi$ (bounded) $1/\varphi$ (but $\nu^{\text{per}}$ bounded)
Net efficiency $\nu(W)$ $\to+\infty$ const.
Coding entropy $H(P)$ $H^{\text{per}}$ ($m$-LRPD sources) $\geq h_{F_{m^{*}}}(P)$
Redundancy $R_{m}(P)$ $O(e^{-\varphi^{m}/\lambda})$ (super-exp.) $\Omega(e^{-F_{m^{*}}/\lambda})$ (fixed)
Dict. value $\mathcal{V}$ $\Omega(\varphi^{k_{\text{max}}})$ (exp. growth) $O(\varphi^{m^{*}})$ (const.)
The aperiodic advantage is a structural consequence of three mathematical properties:
  1. 1.
    $\varphi$ is a Pisot-Vijayaraghavan number (eigenvalue separation and unique recognisability);
  2. 2.
    the Fibonacci word is Sturmian (balance and minimal factor complexity);
  3. 3.
    Weyl’s equidistribution holds at all hierarchy scales.
No periodic tiling possesses any of these three properties.
The Aperiodic Hierarchy Advantage Theorem (Theorem 1) collects these properties into a single characterisation. Table 4 gives the quantitative comparison; Figure 4 and the ablation study (Section 7) provide empirical confirmation.
5. Experimental Setup
5.1 Benchmarks
[…]
  • alice29.txt (Canterbury Corpus [26]): 152,089 bytes, 36,395 word tokens. Classic compression benchmark.
  • enwik8_3M: First 3,000,000 bytes of enwik8 (820,930 word tokens).
  • enwik8_10M: First 10,000,000 bytes of enwik8 (2,751,434 word tokens).
  • enwik8 (Large Text Compression Benchmark [25]): 100,000,000 bytes of Wikipedia XML, 27,731,124 word tokens.
  • enwik9: 1,000,000,000 bytes of Wikipedia XML, 298,263,298 word tokens.
5.2 Baselines
We compare against:
  • gzip -9: zlib DEFLATE, maximum compression level.
  • bzip2 -9: BWT-based, maximum compression.
  • xz -9: LZMA2, maximum compression.
  • All-unigram QTC: A/B test variant with every tile forced to $S$ (no bigram or n-gram lookups), using identical codebooks and escape streams. This isolates the QC contribution.
  • Period-5 QTC: A/B test variant using the LLSLS periodic tiling (same L/S ratio as Fibonacci), with full hierarchy attempted. This isolates the aperiodic advantage.
5.3 Implementation
Quasicryth v5.5 is implemented in ANSI C99. The implementation uses 24-bit arithmetic coding with range coding, bzip2 via libbz2, and zlib for codebook compression. No SIMD or hardware-specific optimisations are used. The codebase is approximately 4,000 lines of C. All experiments were run on a single core; no multi-threading is employed.
5.4 A/B Test Design
The A/B test is a controlled experiment that evaluates the payload contribution of three parsing strategies, using identical codebooks (built once from the input), identical escape streams (all three strategies escape the same OOV words), and identical arithmetic coding models. Only the payload bytes differ. This design ensures that any difference in payload size is due solely to the parsing strategy, not to codebook or escape differences.
6. Results
6.1 Compression Performance
Table 5 reports compression ratios and sizes. On enwik9, Quasicryth achieves 35.99% (359,883,431 B), with the full breakdown in Table 6.
Table 5: Compression ratios (compressed/original $\times 100\%$) for Quasicryth v5.5 versus standard compressors.
Compressor alice29 3MB 10MB enwik8 enwik9
Quasicryth v5.5 36.92 39.29 38.51 37.70 35.99
gzip -9 35.63 36.29 36.85 36.45 32.26
bzip2 -9 28.41 28.89 29.16 29.01 25.40
xz -9 31.89 28.39 27.21 24.87 21.57
Table 6: Compressed file breakdown for enwik9 (1 GB). All sizes in bytes.
Component Size (B)
Original 1,000,000,000
Payload (AC) 254,952,735
Escapes (bz2) 83,537,358
Codebook (zlib) 77,585
Case flags (AC) 21,315,698
Total 359,883,431
Ratio 35.99%
6.2 Deep Hierarchy Hit Counts
Table 7 shows the number of deep hierarchy hits (13-gram and above) across all test files. These positions are exclusively available to the Fibonacci tiling; Period-5’s hierarchy collapses at level 4, providing zero hits at levels 4–9.
Table 7: Deep hierarchy hits by level and file size. All counts are from the Fibonacci tiling; Period-5 achieves zero at every level shown (hierarchy collapsed). Levels 8–9 (89g, 144g) first activate at enwik9 scale.
File Words 13g 21g 34g 55g 89g 144g Total deep
alice29.txt 36K 2 0 0 0 0 0 2
enwik8_3M 821K 971 195 77 0 0 0 1,243
enwik8_10M 2.8M 2,715 383 140 47 0 0 3,285
enwik8 27.7M 24,910 3,256 1,369 551 0 0 30,086
enwik9 298.3M 652,124 109,492 40,475 7,222 2,396 945 812,654
[…]
Figure 4: Deep hierarchy hits (levels 4–9, Fibonacci-exclusive) per word versus corpus size (log-log). Labels show absolute hit counts. The ratio is roughly flat from enwik8_3M to enwik8 while fixed-size codebooks limit density; at enwik9, levels 8–9 activate and the ratio jumps to $2.7\times 10^{-3}$, confirming new $O(W)$ terms.
152 KB3 MB10 MB100 MB1 GB0%5050%100100%Word-weighted share of deep hits13-gram21-gram34-gram55-gram89-gram144-gram
[…]
6.3 Timing
[…]
Table 8: Compression and decompression times (C implementation, single core).
File Size Compress Decomp. C/D
alice29.txt 152 KB 0.1s <<0.1s
[…]
6.4 Asymmetric Compression Profile
Quasicryth is an inherently asymmetric compressor: the computational cost is concentrated entirely in compression, while decompression is fast and lightweight. This asymmetry is a direct structural consequence of the algorithm, not an optimisation artifact, and it makes Quasicryth particularly suited to write-once, read-many scenarios such as archival storage, content distribution, and static web assets.
Why compression is expensive. Compression performs three costly operations: (1) the phase search evaluates 32 candidate tilings with the full substitution hierarchy, scoring each against the codebooks; (2) codebook construction counts all n-gram frequencies up to 144 words per entry, with periodic pruning passes for the 89-gram and 144-gram levels at the 10 M+ word scale; (3) the 32-phase scoring loop must build the complete hierarchy at each candidate phase, repeating the $O(W)$ hierarchy detection.
Why decompression is cheap. The decompressor reads the 2-byte phase from the header and regenerates the entire tiling deterministically — no search, no n-gram counting, no frequency tables. It then decompresses the zlib’d codebook (77 KB at enwik9 scale), decompresses the bz2 escape buffer, and streams through the AC payload once, consulting preloaded codebook hash tables. The entire decompression is a single sequential pass.
Growing asymmetry at scale. The C/D ratio increases from $6.3\times$ at 3 MB to $15.9\times$ at 1 GB (Figure 6), because the 89-gram and 144-gram frequency counting — which hashes windows of 89–144 consecutive word IDs — dominates codebook construction at the enwik9 scale. Decompression throughput remains roughly constant at $\sim$10 MB/s regardless of file size, while compression throughput falls to $\sim$0.65 MB/s at 1 GB.
Practical implication. For any file decompressed $k$ times, the amortised cost per access is $T_{C}/k+T_{D}$. At enwik9 scale ($T_{C}=1{,}540$ s, $T_{D}=97$ s), even a single decompression amortises the compression cost within $\sim$16 accesses. The upfront compression investment is a one-time payment; every subsequent decompression is fast.
152 KB3 MB10 MB100 MB1 GB055101015152020556.36.311.511.514.214.215.915.9Compress / Decompress ratio
Figure 6: Compression-to-decompression time ratio (C/D ratio) versus file size. The ratio grows from $6.3\times$ at 3 MB to $15.9\times$ at 1 GB as the 89-gram and 144-gram frequency counting begins to dominate codebook construction. Decompression throughput remains stable at $\sim$10 MB/s across all scales; the growing asymmetry is entirely a compression-side effect.
7. Ablation Study
7.1 QC Contribution vs All-Unigram Baseline
Table 9 reports the payload savings of the Fibonacci tiling over an all-unigram baseline (every tile is $S$, no bigram or n-gram attempts), using identical codebooks and escape streams. This isolates the compression value of the entire multi-level parsing structure.
Table 9: Quasicrystal contribution: payload savings over all-unigram baseline.
File Size Ratio QC saves Savings %
alice29.txt 152 KB 36.92% 1,276 B 0.84%
enwik8_3M 3 MB 39.29% 94,321 B 3.14%
enwik8_10M 10 MB 38.51% 312,638 B 3.13%
enwik8 100 MB 37.70% 2,838,937 B 2.84%
enwik9 1,000 MB 35.99% 45,608,715 B 4.56%
The QC contribution increases at larger scales, reaching 4.56% of the original file size at enwik9. This growing advantage reflects the activation of deep hierarchy levels with increasing input size.
7.2 Aperiodic Advantage: Fibonacci vs Period-5
Table 10 reports the A/B payload comparison. Period-5 uses the LLSLS periodic tiling with the same L/S ratio as Fibonacci; both use identical codebooks and escape streams.
Table 10: A/B test: payload-only comparison (bytes). Fibonacci vs Period-5 and All-Unigram, identical codebooks and escapes.
File All-Unigram Period-5 Fibonacci
enwik8_3M 895,581 802,632 801,260
enwik8_10M 2,966,773 2,659,942 2,654,135
enwik8 29,036,210 26,237,097 26,196,712
enwik9 300,561,450 256,302,106 254,952,735
Table 11: Aperiodic advantage: Fibonacci over Period-5 (bytes saved in payload).
File Fib saves over P5 P5 saves over Uni
enwik8_3M 1,372 B 92,949 B
enwik8_10M 5,807 B 306,831 B
enwik8 40,385 B 2,799,113 B
enwik9 1,349,371 B 44,259,344 B
[…]
152 KB3 MB10 MB100 MB1 GB0%22%44%66%File size (MB)QC savings (% of original)
Figure 7: QC tiling contribution (payload savings over all-unigram baseline, as % of original) versus file size, semi-log scale. The contribution rises to 4.56% at 1 GB as deeper hierarchy levels progressively activate.
[…]
7.3 Ablation: Version Progression
Table 12 shows the improvement across versions, using alice29.txt as the reference. Each version adds one architectural element; the dominant gains come from moving to word-level parsing (v5.1) and adding the bz2 escape stream with trigram support (v5.2).
Table 12: Version progression on alice29.txt (152,089 B). “QC saves” is the payload advantage of Fibonacci over all-unigram.
Version Ratio QC saves Key change
v5.0 (byte) 49.4% 394 B QC as context only
v5.1 (word) 43.5% 1,254 B LL=bigram, SS=unigram
v5.2 (multi) 36.0% 1,579 B +trigrams, bz2 escape
v5.3 (deep) 36.9% 1,096 B +5g/8g/13g/21g
v5.4 (full) 36.92% 1,276 B +34g/55g (8 levels)
v5.5 (deep Fib) 36.92% 1,276 B +89g/144g (10 levels)
The 89-gram and 144-gram levels (v5.5) do not improve alice29.txt compression — at 36K words there are insufficient repetitions to populate these codebooks. Their contribution is visible only at the enwik9 scale.
7.4 Approaches Rejected
The following design alternatives were tested and abandoned:
  • Word-level LZ pre-pass. Finding and encoding repeated word sequences as back-references improved overall compression but destroyed the QC contribution (reduced to 334 B on alice29) because the LZ layer captures the same long-range redundancy exploited by the bigram/n-gram codebooks; the two mechanisms compete.
  • Adaptive online codebook (LZW-style). Eliminated static codebook overhead but cold-start penalty outweighed savings on medium-sized inputs; QC contribution became negative on small files.
  • Previous-tile-index context. Using the previous codebook index as an additional context bin created 128 sub-models, diluting training data excessively.
  • All-L tiling. Forcing all tiles to $L$ (always bigram) outperforms Fibonacci at level 0 because every L tile is at least as good as an $S$ tile. However, all-L has no hierarchy (every tile is the same type), yielding zero trigram and n-gram positions. The multi-level hierarchy is where Fibonacci’s advantage lies.
  • More hierarchy context bits. Adding more context bits from the substitution structure helped periodic tilings more than Fibonacci, because Period-5’s regular hierarchy creates more consistent patterns for the statistical models. The aperiodic advantage comes from deep hierarchy positions, not from context diversity.
8. Conclusion
We have presented Quasicryth v5.5, a lossless text compressor based on a deep Fibonacci quasicrystal substitution hierarchy, and proved that this hierarchy never collapses while all periodic alternatives do. The proof synthesises Perron-Frobenius analysis of the substitution matrix, the Pisot-Vijayaraghavan property of $\varphi$, Sturmian balance and aperiodicity, and Weyl’s equidistribution theorem. These are not independent observations: they are aspects of a single mathematical object — the Fibonacci quasicrystal — that jointly guarantee the compression properties of Quasicryth.
The aperiodic advantage is proven and measurable: 1,349,371 B at enwik9 (1 GB), growing superlinearly because new hierarchy levels (89-gram, 144-gram) activate at scale. At enwik9, 812,654 deep hierarchy positions (levels 4–9) are exploited; Period-5, the best periodic approximant with the same L/S ratio, achieves zero deep positions, because its hierarchy collapses at level 4 — exactly as the theory predicts.
To our knowledge, Quasicryth is the first compressor in which the aperiodicity of the parsing structure provides a provable, structurally motivated, and empirically quantified advantage over all periodic alternatives.
8.1 Future Work
  • Even deeper hierarchy. Levels 10–11 (233-gram, 377-gram) would activate at approximately 3 GB and 30 GB respectively. The hierarchy never collapses; each new level adds $O(W)$ to the aperiodic advantage.
  • Comprehensive benchmarking. Systematic evaluation on Canterbury, Silesia, and Calgary corpora.
  • Higher-order context mixing. PAQ-style context mixing where the substitution hierarchy provides structured context at multiple scales, with online-learned blend weights.
  • Two-dimensional extensions. For image data, Penrose or Ammann-Beenker tilings could provide hierarchical context with richer tile-type alphabets ($\geq 4$ tile types), extending the quasicrystalline compression principle to 2D.